a train moving at a constant speed travels 180 miles in 4 hours. how far does the train travel in 7 hours?
Answer by stanbon(75887)

a train moving at a constant speed travels 180 miles in 4 hours. how far does the train travel in 7 hours?
---
time and distance are directly related so
you can use a proportion to solve.
----
Proportion:
d/7 = 180/4
d = 7(45)
d = 315 miles
================
Cheers,
Stan H.